OpenStudy (anonymous):

Simplify and leave your answer in factored form. I'm going to post it after this! (:

OpenStudy (anonymous):

\[ \frac{ \frac{ (x-5) }{ (x+5) } + \frac{ (x-1) }{ (x+1) } }{ \frac{ x }{ (x+1) }-\frac{ (3x-7) }{ x } }\]

OpenStudy (anonymous):

In the end I got (2x(x^2 - 5)) / ((x +5)(2x^2-4x-7))

jimthompson5910 (jim_thompson5910):

Multiply EVERY term by the inner LCD x(x+1)(x+5) So x(x+1)(x+5)[ (x-5)/(x+5)] = x(x+1)(x-5) ------------- x(x+1)(x+5)[ (x-1)/(x+1) ] = x(x+5)(x-1) -------------- x(x+1)(x+5)[ x/(x+1) ] = x^2(x+5) -------------- x(x+1)(x+5)[ (3x-7)/x ] = (x+1)(x+5)(3x-7) =========================================== This means that \[ \frac{ \frac{ (x-5) }{ (x+5) } + \frac{ (x-1) }{ (x+1) } }{ \frac{ x }{ (x+1) }-\frac{ (3x-7) }{ x } }\] turns into \[ \frac{ x(x+1)(x-5) + x(x+5)(x-1) }{ x^2(x+5)-(x+1)(x+5)(3x-7) }\] I'll let you finish this

OpenStudy (anonymous):

Oh, thanks!

jimthompson5910 (jim_thompson5910):

Let me know what you get for the final answer